OpenStudy (epicteatime):

Justify if completing the square is a good method for solving when the Discriminant is negative. Use any of your three functions as an example and respond in complete sentences. My beautiful function: x2-10x+24

OpenStudy (epicteatime):

@amistre64

OpenStudy (epicteatime):

@gorv

OpenStudy (epicteatime):

@jim_thompson5910

OpenStudy (epicteatime):

@mathstudent55

OpenStudy (epicteatime):

HALP WATER

OpenStudy (anonymous):

What I am supposed to do?? Completing the square of that function?

OpenStudy (epicteatime):

i mean, when the discriminant is a negative, doesn't that mean the answer is going be a complex root thing?

OpenStudy (epicteatime):

wait, so what should i do?

OpenStudy (epicteatime):

which formula should i use to make it a "perfect square"?

OpenStudy (anonymous):

When D is negative, you are surely going to get Complex Roots, do you have any doubt?

OpenStudy (epicteatime):

nope

OpenStudy (epicteatime):

what else should i write...

OpenStudy (anonymous):

Are you trying to find x or doing what? What exactly you are doing?

OpenStudy (epicteatime):

justify if completing the square of that equation is good when your trying to get answer with a negative discriminant. Well i technically just "think" thats what it means, i came here to see if anyone could "understand" the question.

OpenStudy (epicteatime):

cause i have no idea what the hell that thing is saying.

OpenStudy (anonymous):

Same here, but let us start completion of square method, do you know it?

OpenStudy (epicteatime):

somewhat

OpenStudy (epicteatime):

i keep forgetting that formula

OpenStudy (anonymous):

That will do, don't worry.. :)

OpenStudy (epicteatime):

was it the one with a^2(y+h)-f or something?

OpenStudy (epicteatime):

or was it the other other one?

OpenStudy (anonymous):

It is just: \[(a+b)^2 = a^2 + 2ab + b^2 \\ (a-b)^2 = a^2 -2ab + b^2\]

OpenStudy (epicteatime):

:O

OpenStudy (anonymous):

Just basic formulae.. :)

OpenStudy (epicteatime):

so this is useless? (x-4)(x-6), x2-10x+24, b2-4ac, -102-4(1)(24), 100-96, 4

OpenStudy (anonymous):

\(x^2 - 10x + 24\) \(D = 100 - 96 = 4\) D is not negative.. :(

OpenStudy (epicteatime):

oh damn

OpenStudy (epicteatime):

WE SHALL USE ThE OTHER OTHER ONE

OpenStudy (anonymous):

Choose function with D = negative..

OpenStudy (epicteatime):

x2-10x+28, b2-4ac: -102-4(1)(28), 100-112, -12.

OpenStudy (epicteatime):

ther

OpenStudy (epicteatime):

i made 3 extras XD

OpenStudy (epicteatime):

the discriminant is -12

OpenStudy (epicteatime):

yay...

OpenStudy (anonymous):

Let amistre keep his point here, we should not do without his guidance.. :P

OpenStudy (epicteatime):

huzzah!

OpenStudy (amistre64):

quadratic formula is a formula that is developed from completing the square .... long way, complete the square; short way: use the formula

OpenStudy (epicteatime):

soooooo

OpenStudy (anonymous):

Yeah I never thought it that way.. :(

OpenStudy (amistre64):

so, we prove the formula by completing the square ....

OpenStudy (anonymous):

Yep..

OpenStudy (epicteatime):

yeah...

OpenStudy (epicteatime):

wait so how do i input the information into the formula water gave me?

OpenStudy (anonymous):

That is why I asked you if you are finding x then we have two methods to use.. We can use Quadratic Formula as well as Factorization if applicable.. :)

OpenStudy (epicteatime):

quadratic formula pls

OpenStudy (anonymous):

Yes, you draw nice.. :P

OpenStudy (epicteatime):

thank you

OpenStudy (epicteatime):

i think amistre is writing a very long thingy

OpenStudy (amistre64):

\[ax^2 + bx + c = 0\] \[ax^2 + bx =-c\] \[x^2 + \frac bax =-\frac ca\] \[x^2 + \frac bax+\frac{b^2}{(2a)^2} =\frac{b^2}{(2a)^2}-\frac ca\] \[(x+ \frac ba)^2 =\frac{b^2}{(2a)^2}-\frac ca\] \[x+ \frac ba =\pm\sqrt{\frac{b^2}{(2a)^2}-\frac ca}\] \[x=- \frac ba \pm\sqrt{\frac{b^2}{(2a)^2}-\frac ca}\]

OpenStudy (anonymous):

\[D = \sqrt{b^2-4ac} \\ x = \frac{-b \pm D}{2a}\]

OpenStudy (epicteatime):

oh that formula

OpenStudy (epicteatime):

moving the number to the other side

OpenStudy (amistre64):

ugh, i dropped a 2

OpenStudy (amistre64):

(x+b/(2a))^2 and work it thru

OpenStudy (epicteatime):

x2-10x+28, x2-10x=-28, (x2-10x+25)=-28+25, (x2-10x+25)=3, (x-5)2=-3,

OpenStudy (epicteatime):

is it that?

OpenStudy (epicteatime):

then you root it

OpenStudy (epicteatime):

\[\sqrt{x-5}^{2}=\sqrt{3}\]

OpenStudy (amistre64):

\[x=- \frac b{2a} \pm\sqrt{\frac{b^2}{(2a)^2}-\frac ca}\] \[x=- \frac b{2a} \pm\sqrt{\frac{b^2}{4a^2}-\frac ca}\] \[x=- \frac b{2a} \pm\sqrt{\frac{b^2}{4a^2}-\frac {4ac}{4a^2}}\] \[x=- \frac b{2a} \pm\frac{\sqrt{b^2-4ac}}{2a}\]

OpenStudy (epicteatime):

it seems your doing it another way....

OpenStudy (amistre64):

im proving the formula is all by taking a general quadratic and algebraing it into the formula .... solving for x

OpenStudy (epicteatime):

i don't think i learnt that...

OpenStudy (amistre64):

and D is technically equal to the inside of the sqrt, not the whole sqrt .... but i prefer to include the sqrt as water did

OpenStudy (amistre64):

you didnt learn how to complete a square?

OpenStudy (epicteatime):

I'm just confused tis all

OpenStudy (amistre64):

did you cover how to complete a square?

OpenStudy (epicteatime):

i think

OpenStudy (epicteatime):

wait I'm on the lesson thats called "completing a square"

OpenStudy (epicteatime):

well #### this

OpenStudy (amistre64):

the method of completing the square, results in the quadratic formula .. it is the proof.

OpenStudy (epicteatime):

its just i never saw that "formula" while i tried to stay awake in class

OpenStudy (epicteatime):

hmmm, its either i seriously am being derby, or its just that i didn't learn it

OpenStudy (epicteatime):

i only learnt the thing i wrote

OpenStudy (epicteatime):

and then the +- 5 and 3

OpenStudy (epicteatime):

and then you figure both of those out

OpenStudy (amistre64):

given: ax^2 + bx + c = 0 what is the first step in completing the square?

OpenStudy (epicteatime):

you move the c over to 0

OpenStudy (epicteatime):

which makes it -c

OpenStudy (amistre64):

ax^2 + bx = -c then?

OpenStudy (epicteatime):

then you divide bx by 2, and square that

OpenStudy (amistre64):

nope, need to get rid of that a

OpenStudy (epicteatime):

oh

OpenStudy (amistre64):

divide off the a

OpenStudy (epicteatime):

if a has a number other than 1

OpenStudy (epicteatime):

than you factor it

OpenStudy (epicteatime):

into a(x+bx)=-c

OpenStudy (amistre64):

a=1 still works when you divide by it :) x^2 + b/a x = -c/a now we can do the halving thing

OpenStudy (epicteatime):

x2

OpenStudy (epicteatime):

yeah....

OpenStudy (amistre64):

do you see that we divided both sides by a?

OpenStudy (epicteatime):

yep

OpenStudy (amistre64):

ax^2 + bx = -c x^2 + b/a x = -c/a ok, now lets add the (b/2a)^2 to each side x^2 + b/a x + (b/2a)^2 = (b/2a)^2 - c/a whats next?

OpenStudy (epicteatime):

uhh

OpenStudy (epicteatime):

combining them all?

OpenStudy (epicteatime):

or like terms

OpenStudy (epicteatime):

cause i have definitely NOT learnt this

OpenStudy (amistre64):

turning the left side into a (x+b/2a)^2 yes since it is not a perfect square on that side